OpenStudy (anonymous):

find the sum in simplest form. 2-g/g^6 - 1/g^6

OpenStudy (anonymous):

\[\frac{ 2-g }{ g^6 }-\frac{ 1 }{ g^6 }\]

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

you have same denominators, just subtract the numenators

OpenStudy (anonymous):

2-g-1=1-g

OpenStudy (anonymous):

1-g/g^6

OpenStudy (anonymous):

thank you